1. Anne and Nancy use a metal mixture that is 20% copper to make jewelry. How many ounces of a 14% alloy must be mixed with a 23% alloy to form 69 ounces of the desired alloy?

2. A cruise boat travels 36 miles downstream in 3 hours and returns to its starting point in 12 hours. Find the speed of the stream?

Answer by stanbon(75887)

1. Anne and Nancy use a metal mixture that is 20% copper to make jewelry. How many ounces of a 14% alloy must be mixed with a 23% alloy to form 69 ounces of the desired alloy?
----
copper + copper = copper
0.14x + 0.23(69-x) = 0.20*69
14x + 23*69 - 23x = 20*69
-9x = -3*69
x = 23 oz (amt. of 14% needed)
69-x = 46 oz (amt. of 23% needed)
------------
2. A cruise boat travels 36 miles downstream in 3 hours and returns to its starting point in 12 hours. Find the speed of the stream?
downstream DATA:
dist = 36 miles ; time = 3 hrs ; rate = 36/3 = 12 mph
----
upstream DATA:
dist = 36 miles ; time = 12 hrs ; rate = 36/12 = 3 mph
-----
b + s = 12 mph
b - s = 3 mph
----
2b = 15 mph
boat speed = 7.5 mph
---
s = 12-b = 4.5 mph (stream speed)
------------------
Cheers,
Stan H.
--------------